# High-dimensional sliced inverse regression with endogeneity

### v1

@@ Acknowledgments
$\left\lVert\widehat{\bm{\Omega}}-\bm{\Omega}\right\rVert_{\max}=\left\lVert(\bm{\Sigma}_{\mathbf{Z}}-\widehat{\bm{\Sigma}}_{\mathbf{Z}})+(\mathbf{\Xi}-\widehat{\mathbf{\Xi}})\right\rVert_{\max}\leq\left\lVert\widehat{\bm{\Sigma}}_{\mathbf{Z}}-\bm{\Sigma}_{\mathbf{Z}}\right\rVert_{\max}+\left\lVert\widehat{\mathbf{\Xi}}-\mathbf{\Xi}\right\rVert_{\max}.$
@@
with probability at least $1-\exp(-C^{\prime}\log q)$ (Ravikumar et al.,, 2011; Tan et al.,, 2018). Next, we will bound the elementwise maximum norm $\left\lVert\widehat{\mathbf{\Xi}}-\mathbf{\Xi}\right\rVert_{\max}.$
Following Zhu et al., (2006), we introduce the following notations. Let $y_{(1)},\ldots,y_{(n)}$ denote the order statistics of the outcome data $y_{1},\ldots,y_{n}$, and ${\mathbf{Z}}_{(1)},\ldots,\mathbf{Z}_{(n)}$ be the concomitants in the $\mathcal{Z}$ data. Define $\bm{m}(y_{(i)})=E(\mathbf{Z}_{(i)}\mid y_{(i)})\in\mathbb{R}^{q}$ and $\bm{\epsilon}_{(i)}=\mathbf{Z}_{(i)}-\bm{m}(y_{(i)})$, and let $\mathbf{M}_{n}$ be the $n\times p$ matrix whose $i$th row is $\bm{m}(y_{(i)})$ and $\mathbf{E}_{n}$ be the $n\times p$ matrix whose $i$th row is $\bm{\epsilon}_{i}$. Note we have $E(\bm{\epsilon}_{(i)})=0$ and $\operatorname{\text{Cov}}\left\{\bm{m}(y_{(i)}),\bm{\epsilon}_{(i)}\right\}=\bm{0}$ for all $i=1,\ldots,n$.
Let $Z_{(i)j}$, $m_{j}(y_{(i)})$ and $\epsilon_{(i)j}$ be the $j$th element of $\mathbf{Z}_{(i)}$, $\bm{m}(y_{(i)})$ and $\epsilon_{(i)}$, respectively, for $j=1,\ldots,q$. Similar to Proposition 2 in the Supplementary Materials of Tan et al., (2018), we obtain the following properties of $\bm{m}(y_{i})$ and $\epsilon_{(i)}$, whose proof is omitted.
@@
Under Condition 4.1, $Z_{(i)j}$, $m_{j}(y_{(i)})$ and $\epsilon_{(i)j}$ are sub-Gaussian with sub-Gaussian norms $\left\lVert\mathbf{Z}_{j}\right\rVert_{\psi_{2}}\leq C\sqrt{(\bm{\Sigma}_{\mathbf{Z}})_{jj}}$, $\left\lVert\mathbf{Z}_{j}\right\rVert_{\psi_{2}}$ and $2\left\lVert\mathbf{Z}_{j}\right\rVert_{\psi_{2}}$, respectively, for $i=1,\ldots,n,~{}j=1,\ldots,q$.
@@
$\widehat{\mathbf{\Xi}}$ $=\frac{1}{H(c-1)}\left\{\left(1-\frac{1}{c}\right)\mathbf{E}_{n}^{\top}\mathbf{E}_{n}\right\}+\frac{1}{H(c-1)}\left\{\mathbf{E}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{Q}_{c}\right)\mathbf{E}_{n}\right\}$ (33)
@@
$:=\mathbf{J}_{1}+\mathbf{J}_{2}+\mathbf{J}_{3}+\mathbf{J}_{4},$
@@
$\|\mathbf{\Xi}-\widehat{\mathbf{\Xi}}\|_{\max}\leq\|\mathbf{J}_{1}-{\mathbf{\Xi}}\|_{\max}+\left\lVert\mathbf{J}_{2}\right\rVert_{\max}+\left\lVert\mathbf{J}_{3}\right\rVert_{\max}+\left\lVert\mathbf{J}_{4}\right\rVert_{\max}.$ (34)
@@
For the first term $\|\mathbf{J}_{1}-{\mathbf{\Xi}}\|_{\max}$, noting that $\mathbf{J}_{1}=n^{-1}\mathbf{E}_{n}^{\top}\mathbf{E}_{n}=n^{-1}\sum_{i=1}^{n}\bm{\epsilon}_{i}\bm{\epsilon}_{i}^{\top}$, then $\text{E}(\mathbf{J}_{1})=\mathbf{\Xi}$. On combining this with Proposition 5, it follows that each element of $\mathbf{J}_{1}-\mathbf{\Xi}$ is a centered sub-exponential random variable with sub-exponential norm bounded by $CL^{2}$, since $\left\lVert\mathbf{Z}\right\rVert_{\psi_{2}}\leq L$ by condition 1. Therefore, with probability at least $1-\exp(-C^{\prime}\log q)$, we have
$\|\mathbf{J}_{1}-{\mathbf{\Xi}}\|_{\max}\leq C\sqrt{\frac{\log q}{n}}.$ (35)
For term $\left\lVert\mathbf{J}_{2}\right\rVert_{\max}$, we note that $\mathbf{Q}_{c}$ has all diagonal elements zero and off-diagonal elements $-c^{-1}$, so each element of $\mathbf{J}_{2}$ is the sum of $n$ terms of the form $[H(c-1)]^{-1}\epsilon_{ij}\epsilon_{i^{\prime}j^{\prime}}$ with $i\neq i^{\prime}$ (note $j$ and $j^{\prime}$ can be the same). Next, conditional on the order statistics $y_{(1)},\ldots,y_{(n)}$, the quantities $\bm{\epsilon}_{i}$ and $\bm{\epsilon}_{i^{\prime}}$ are independent (Yang,, 1977). On combining this with Proposition 5, we have that $\epsilon_{ij}\epsilon_{i^{\prime}j^{\prime}}$ is zero-mean sub-exponential with sub-exponential norm bounded by $CL^{2}$. Therefore, with probability at least $1-\exp(-C^{\prime}\log q)$, we have
@@
Next, we consider the term $\mathbf{J}_{3}$, where it suffices to consider the term $\mathbf{E}_{n}^{\top}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})\mathbf{M}_{n}$. Let $\widetilde{\mathbf{M}}_{n}$ denote the $n\times p$ matrix whose $i$th row is given by $\bm{m}(y_{(i)})-\bm{m}(y_{(i+1)})$ for $11\leq i\leq n-$ 1, and whose $n$th row is $\bm{y}_{(n)}$. It is straightforward to see that $\widetilde{\mathbf{M}}_{n}=\mathbf{A}_{n}\mathbf{M}_{n}$ and $\mathbf{M}_{n}=\mathbf{A}_{n}^{-1}\widetilde{\mathbf{M}}_{n}$, where for any $t$, we have
@@
$\left\lVert\mathbf{E}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{P}_{c}\right)\mathbf{M}_{n}\right\rVert_{\max}$ $=\left\lVert\mathbf{E}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{P}_{c}\right)\mathbf{A}_{n}^{-1}\widetilde{\mathbf{M}}_{n}^{T}\right\rVert_{\max}$
$=\left\lVert\mathbf{E}_{n}^{\top}\left\{\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}^{-1}\right)\right\}\widetilde{\mathbf{M}}_{n}\right\rVert_{\max}$
$\leq\left\lVert\mathbf{E}_{n}\right\rVert_{\max}\left\lVert\left\{\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}^{-1}\right)\right\}\widetilde{\mathbf{M}}_{n}\right\rVert_{1}.$
We have $\left\lVert\mathbf{E}_{n}\right\rVert_{\max}=\max_{i,j}|\epsilon_{ij}|$, where all $\epsilon_{i,j}$ are sub-Gaussian random variables by Proposition 5 for all $i=1,\ldots,n$ and $j=1,\ldots,q$. Therefore, with probability at least $1-\exp(C^{\prime}\log q)$, we have that $\left\lVert\mathbf{E}_{n}\right\rVert_{\max}\leq C\sqrt{\log d}$. Also note that in the matrix $\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)$, all the columns of $\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)$ have their sum of absolute values less than two, i.e $\left\lVert\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)\right\rVert_{1}\leq 2$. On combining these results, we obtain
$\left\lVert\left\{\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)\right\}\widetilde{\mathbf{M}}_{n}\right\rVert_{1}\leq 2\left\lVert\widetilde{\mathbf{M}}_{n}\right\rVert_{1}\leq\sum_{i=2}^{n}\left\lVert\bm{m}\left(y_{(i)}\right)-\bm{m}\left(y_{(i-1)}\right)\right\rVert_{\infty}\leq Cn^{1/4},$
@@
$\left\lVert\mathbf{J}_{3}\right\rVert_{\max}\leq C\left(\dfrac{1}{n}\sqrt{\log q}~{}n^{1/4}\right)=\dfrac{C}{n^{1/4}}\sqrt{\frac{\log q}{n}}.$ (37)
@@
$\left\lVert\mathbf{M}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{P}_{c}\right)\mathbf{M}_{n}\right\rVert_{\max}=\left\lVert\widetilde{\mathbf{M}}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{A}_{c}^{-\top}\mathbf{P}_{c}\mathbf{A}_{c}^{-1}\right)\widetilde{\mathbf{M}}_{n}\right\rVert_{\max}.$
@@
$\left\lVert\mathbf{J}_{4}\right\rVert_{\max}\leq\dfrac{C}{n}\left\lVert\widetilde{\mathbf{M}}_{n}\right\rVert_{1}^{2}\leq\dfrac{C}{n}\left[\sup_{\Pi_{n}(B)}\sum_{i=2}^{n}\left\|m\left\{y_{(i)}\right\}-m\left\{y_{(i-1)}\right\}\right\|_{\infty}\right]^{2}\leq\dfrac{C}{\sqrt{n}},$ (38)
@@
$\left\lVert\mathbf{\Xi}-\widehat{\mathbf{\Xi}}\right\rVert_{\max}\leq C\sqrt{\frac{\log q}{n}},$ (39)
@@
$\left\lVert\widehat{\bm{\Omega}}-\bm{\Omega}\right\rVert_{\max}\leq C\sqrt{\frac{\log q}{n}},$
@@
$\left\lVert\widehat{\bm{\Sigma}}^{1/2}_{\widehat{\mathbf{X}}}\bm{\delta}\right\rVert_{2}^{2}=\bm{\delta}^{\top}\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\bm{\delta}=\bm{\delta}^{\top}(\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}})\bm{\delta}+\bm{\delta}^{\top}\bm{\Sigma}_{\widehat{\mathbf{X}}}\bm{\delta}$ $\geq\kappa_{1}\left\lVert\bm{\delta}\right\rVert_{2}^{2}-\bm{\delta}^{\top}(\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}})\bm{\delta},$ (40)
@@
$\bm{\delta}^{\top}(\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}})\bm{\delta}\leq\left\lVert\bm{\delta}\right\rVert_{1}^{2}\left\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\right\rVert_{\max}.$
Since $\bm{\delta}\in\mathcal{C}(S,3)$, i.e $\left\lVert\bm{\delta}_{S^{c}}\right\rVert_{1}\leq 3\left\lVert\bm{\delta}_{S}\right\rVert_{1}$, then we obtain $\left\lVert\bm{\delta}\right\rVert_{1}^{2}=(\left\lVert\bm{\delta}_{S}\right\rVert_{1}+\left\lVert\bm{\delta}_{S^{c}}\right\rVert_{1})^{2}\leq 16\left\lVert\bm{\delta}_{S}\right\rVert_{1}^{2}\leq 16s\left\lVert\bm{\delta}\right\rVert_{2}^{2}$. Furthermore, noting that $\bm{\Sigma}_{\widehat{\mathbf{X}}}=\bm{\Gamma}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\Gamma}$ and $\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}=\widehat{\bm{\Gamma}}^{\top}\widehat{\bm{\Sigma}}_{\mathbf{Z}}\widehat{\bm{\Gamma}}$, then we have
$\left\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\right\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\widehat{\bm{\Sigma}}_{\mathbf{Z}}\hat{\bm{\gamma}}_{k}|.$
@@
The three terms $\tilde{T}_{1}$, $\tilde{T}_{2}$ and $\tilde{T}_{3}$ are similar to $T_{1},T_{2}$ and $T_{3}$ in Section S2.5 with $\bm{\Omega}$ and $\widehat{\bm{\Omega}}$ replaced by $\bm{\Sigma}_{\mathbf{Z}}$ and $\widehat{\bm{\Sigma}}_{\mathbf{Z}}$, respectively. It follows that by a similar argument, we obtain
$\left\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\right\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\bm{\Sigma}_{\widehat{\mathbf{X}}}\hat{\bm{\gamma}}_{k}|\leq CL\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$ (42)
@@
$\left\lVert\widehat{\bm{\Sigma}}^{1/2}_{\widehat{\mathbf{X}}}\bm{\delta}\right\rVert_{2}^{2}\geq\kappa_{2}\left\lVert\bm{\delta}\right\rVert_{2}^{2},$

### v2

@@ C.1 Proof of Lemma 3
$\lVert\widehat{\bm{\Omega}}-\bm{\Omega}\rVert_{\max}=\lVert(\bm{\Sigma}_{\mathbf{Z}}-\widehat{\bm{\Sigma}}_{\mathbf{Z}})+(\bm{\Xi}-\widehat{\bm{\Xi}})\rVert_{\max}\leq\lVert\widehat{\bm{\Sigma}}_{\mathbf{Z}}-\bm{\Sigma}_{\mathbf{Z}}\rVert_{\max}+\lVert\widehat{\bm{\Xi}}-\bm{\Xi}\rVert_{\max}.$
@@
with probability at least $1-\exp(-C^{\prime}\log q)$ (Ravikumar et al.,, 2011; Tan et al.,, 2018). Next, we will bound the elementwise maximum norm $\lVert\widehat{\bm{\Xi}}-\bm{\Xi}\rVert_{\max}.$
Following Zhu et al., (2006), we introduce the following notations. Let $y_{(1)},\ldots,y_{(n)}$ denote the order statistics of the outcome data $y_{1},\ldots,y_{n}$, and ${\mathbf{Z}}_{(1)},\ldots,\mathbf{Z}_{(n)}$ be the concomitants in the $\mathcal{Z}$ data. Define $m(y_{(i)})=E(\mathbf{Z}_{(i)}\mid y_{(i)})\in\mathbb{R}^{q}$ and $\bm{\epsilon}_{(i)}=\mathbf{Z}_{(i)}-m(y_{(i)})$, and let $\mathbf{M}_{n}$ be the $n\times p$ matrix whose $i$th row is $m(y_{(i)})$ and $\mathbf{E}_{n}$ be the $n\times p$ matrix whose $i$th row is $\bm{\epsilon}_{i}$. Note we have $E(\bm{\epsilon}_{(i)})=0$ and $\operatorname{\text{cov}}\left\{m(y_{(i)}),\bm{\epsilon}_{(i)}\right\}=0$ for all $i=1,\ldots,n$.
Let $Z_{(i)j}$, $m_{j}(y_{(i)})$ and $\epsilon_{(i)j}$ be the $j$th element of $\mathbf{Z}_{(i)}$, $m(y_{(i)})$ and $\epsilon_{(i)}$, respectively, for $j=1,\ldots,q$. Similar to Proposition 2 in the Supplementary Materials of Tan et al., (2018), we obtain the following properties of $m(y_{i})$ and $\epsilon_{(i)}$, whose proof is omitted.
@@
Under Condition 4.1, $Z_{(i)j}$, $m_{j}(y_{(i)})$ and $\epsilon_{(i)j}$ are sub-Gaussian with sub-Gaussian norms $\lVert\mathbf{Z}_{j}\rVert_{\psi_{2}}\leq C\sqrt{(\bm{\Sigma}_{\mathbf{Z}})_{jj}}$, $\lVert\mathbf{Z}_{j}\rVert_{\psi_{2}}$ and $2\lVert\mathbf{Z}_{j}\rVert_{\psi_{2}}$, respectively, for $i=1,\ldots,n,~j=1,\ldots,q$.
@@
$\widehat{\bm{\Xi}}$ $=\frac{1}{H(c-1)}\left\{\left(1-\frac{1}{c}\right)\mathbf{E}_{n}^{\top}\mathbf{E}_{n}\right\}+\frac{1}{H(c-1)}\left\{\mathbf{E}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{Q}_{c}\right)\mathbf{E}_{n}\right\}$ (43)
@@
$=\mathbf{J}_{1}+\mathbf{J}_{2}+\mathbf{J}_{3}+\mathbf{J}_{4},$
@@
$\|\bm{\Xi}-\widehat{\bm{\Xi}}\|_{\max}\leq\|\mathbf{J}_{1}-{\bm{\Xi}}\|_{\max}+\lVert\mathbf{J}_{2}\rVert_{\max}+\lVert\mathbf{J}_{3}\rVert_{\max}+\lVert\mathbf{J}_{4}\rVert_{\max}.$ (44)
@@
For the first term $\|\mathbf{J}_{1}-{\bm{\Xi}}\|_{\max}$, noting that $\mathbf{J}_{1}=n^{-1}\mathbf{E}_{n}^{\top}\mathbf{E}_{n}=n^{-1}\sum_{i=1}^{n}\bm{\epsilon}_{i}\bm{\epsilon}_{i}^{\top}$, then $\text{E}(\mathbf{J}_{1})=\bm{\Xi}$. On combining this with Proposition 6, it follows that each element of $\mathbf{J}_{1}-\bm{\Xi}$ is a centered sub-exponential random variable with sub-exponential norm bounded by $CL^{2}$, since $\lVert\mathbf{Z}\rVert_{\psi_{2}}\leq L$ by condition 1. Therefore, with probability at least $1-\exp(-C^{\prime}\log q)$, we have
$\|\mathbf{J}_{1}-{\bm{\Xi}}\|_{\max}\leq C\sqrt{\frac{\log q}{n}}.$ (45)
For term $\lVert\mathbf{J}_{2}\rVert_{\max}$, we note that $\mathbf{Q}_{c}$ has all diagonal elements zero and off-diagonal elements $c^{-1}$, so each element of $\mathbf{J}_{2}$ is the sum of $n$ terms of the form $H(c-1)]^{-1}\epsilon_{ij}\epsilon_{i^{\prime}j^{\prime}}$ with $i\neq i^{\prime}$ (note $j$ and $j^{\prime}$ can be the same). Next, conditional on the order statistics $y_{(1)},\ldots,y_{(n)}$, the quantities $\bm{\epsilon}_{i}$ and $\bm{\epsilon}_{i^{\prime}}$ are independent (Yang,, 1977). On combining this with Proposition 6, we have that $\epsilon_{ij}\epsilon_{i^{\prime}j^{\prime}}$ is zero-mean sub-exponential with sub-exponential norm bounded by $CL^{2}$. Therefore, with probability at least $1-\exp(-C^{\prime}\log q)$, we have
@@
Next, we consider the term $\mathbf{J}_{3}$, where it suffices to consider the term $\mathbf{E}_{n}^{\top}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})\mathbf{M}_{n}$. Let $\widetilde{\mathbf{M}}_{n}$ denote the $n\times p$ matrix whose $i$th row is given by $m(y_{(i)})-m(y_{(i+1)})$ for $1\leq i\leq n-1$, and whose $n$th row is $\bm{y}_{(n)}$. It is straightforward to see that $\widetilde{\mathbf{M}}_{n}=\mathbf{A}_{n}\mathbf{M}_{n}$ and $\mathbf{M}_{n}=\mathbf{A}_{n}^{-1}\widetilde{\mathbf{M}}_{n}$, where for any $t$, we have
@@
$\lVert\mathbf{E}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{P}_{c}\right)\mathbf{M}_{n}\rVert_{\max}$ $=\lVert\mathbf{E}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{P}_{c}\right)\mathbf{A}_{n}^{-1}\widetilde{\mathbf{M}}_{n}^{T}\rVert_{\max}$
$=\lVert\mathbf{E}_{n}^{\top}\left\{\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}^{-1}\right)\right\}\widetilde{\mathbf{M}}_{n}\rVert_{\max}$
$\leq\lVert\mathbf{E}_{n}\rVert_{\max}\lVert\left\{\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}^{-1}\right)\right\}\widetilde{\mathbf{M}}_{n}\rVert_{1}.$
We have $\lVert\mathbf{E}_{n}\rVert_{\max}=\max_{i,j}|\epsilon_{ij}|$, where all $\epsilon_{i,j}$ are sub-Gaussian random variables by Proposition 6 for all $i=1,\ldots,n$ and $j=1,\ldots,q$. Therefore, with probability at least $1-\exp(C^{\prime}\log q)$, we have that $\lVert\mathbf{E}_{n}\rVert_{\max}\leq C\sqrt{\log d}$. Also note that in the matrix $\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)$, all the columns of $\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)$ have their sum of absolute values less than two, i.e $\lVert\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)\rVert_{1}\leq 2$. On combining these results, we obtain
$\lVert\left\{\mathbf{I}_{H}\otimes\left(\mathbf{P}_{c}\mathbf{A}_{c}\right)\right\}\widetilde{\mathbf{M}}_{n}\rVert_{1}\leq 2\lVert\widetilde{\mathbf{M}}_{n}\rVert_{1}\leq\sum_{i=2}^{n}\lVert m\left(y_{(i)}\right)-m\left(y_{(i-1)}\right)\rVert_{\infty}\leq Cn^{1/4},$
@@
$\lVert\mathbf{J}_{3}\rVert_{\max}\leq C\left(\dfrac{1}{n}\sqrt{\log q}~n^{1/4}\right)=\dfrac{C}{n^{1/4}}\sqrt{\frac{\log q}{n}}.$ (47)
@@
$\lVert\mathbf{M}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{P}_{c}\right)\mathbf{M}_{n}\rVert_{\max}=\lVert\widetilde{\mathbf{M}}_{n}^{\top}\left(\mathbf{I}_{H}\otimes\mathbf{A}_{c}^{-\top}\mathbf{P}_{c}\mathbf{A}_{c}^{-1}\right)\widetilde{\mathbf{M}}_{n}\rVert_{\max}.$
@@
$\lVert\mathbf{J}_{4}\rVert_{\max}\leq\dfrac{C}{n}\lVert\widetilde{\mathbf{M}}_{n}\rVert_{1}^{2}\leq\dfrac{C}{n}\left[\sup_{\Pi_{n}(B)}\sum_{i=2}^{n}\left\|m\left\{y_{(i)}\right\}-m\left\{y_{(i-1)}\right\}\right\|_{\infty}\right]^{2}\leq\dfrac{C}{\sqrt{n}},$ (48)
@@
$\lVert\bm{\Xi}-\widehat{\bm{\Xi}}\rVert_{\max}\leq C\sqrt{\frac{\log q}{n}},$ (49)
@@
$\lVert\widehat{\bm{\Omega}}-\bm{\Omega}\rVert_{\max}\leq C\sqrt{\frac{\log q}{n}},$
@@ C.2 Proof of Proposition 4
$\lVert\widehat{\bm{\Sigma}}^{1/2}_{\widehat{\mathbf{X}}}\bm{\delta}\rVert_{2}^{2}=\bm{\delta}^{\top}\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\bm{\delta}=\bm{\delta}^{\top}(\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}})\bm{\delta}+\bm{\delta}^{\top}\bm{\Sigma}_{\widehat{\mathbf{X}}}\delta$ $\geq\kappa_{1}\lVert\bm{\delta}\rVert_{2}^{2}-\bm{\delta}^{\top}(\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}})\bm{\delta},$ (50)
@@
$\bm{\delta}^{\top}(\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}})\bm{\delta}\leq\lVert\bm{\delta}\rVert_{1}^{2}\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\rVert_{\max}.$
Since $\bm{\delta}\in\mathcal{C}(S,3)$, i.e $\lVert\bm{\delta}_{S^{c}}\rVert_{1}\leq 3\lVert\bm{\delta}_{S}\rVert_{1}$, then we obtain $\lVert\bm{\delta}\rVert_{1}^{2}=(\lVert\bm{\delta}_{S}\rVert_{1}+\lVert\bm{\delta}_{S^{c}}\rVert_{1})^{2}\leq 16\lVert\bm{\delta}_{S}\rVert_{1}^{2}\leq 16s\lVert\bm{\delta}\rVert_{2}^{2}$. Furthermore, noting that $\bm{\Sigma}_{\widehat{\mathbf{X}}}=\bm{\Gamma}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\Gamma}$ and $\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}=\widehat{\bm{\Gamma}}^{\top}\widehat{\bm{\Sigma}}_{\mathbf{Z}}\widehat{\bm{\Gamma}}$, then we have
$\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\widehat{\bm{\Sigma}}_{\mathbf{Z}}\hat{\bm{\gamma}}_{k}|.$
@@
The three terms $\tilde{T}_{1}$, $\tilde{T}_{2}$ and $\tilde{T}_{3}$ are similar to $T_{1},T_{2}$ and $T_{3}$ in Section B.4 with $\bm{\Omega}$ and $\widehat{\bm{\Omega}}$ replaced by $\bm{\Sigma}_{\mathbf{Z}}$ and $\widehat{\bm{\Sigma}}_{\mathbf{Z}}$, respectively. It follows that by a similar argument, we obtain
$\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\bm{\Sigma}_{\widehat{\mathbf{X}}}\hat{\bm{\gamma}}_{k}|\leq CL\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$ (52)
@@
$\lVert\widehat{\bm{\Sigma}}^{1/2}_{\widehat{\mathbf{X}}}\bm{\delta}\rVert_{2}^{2}\geq\kappa_{2}\lVert\bm{\delta}\rVert_{2}^{2},$